OpenStudy (anonymous):

write in slope intercept form identify slope and y-intercept 2x+3y=6

OpenStudy (anonymous):

slope intercept is y=mx+b so rewrite it interms of y 3y=6-2x y=6/3-2x/3 y=3-2x/3